Scenario #2: The mortgage department of the company is selling two model homes that are located on the same block. The square footage, as well as the type of model, determines the cost of these homes.
e) The second home has 200 square feet more than the square footage of the first home. Write an equation that illustrates this situation using x and y to denote home #1 (x) and home #2 (y), respectively.
f) Together, both homes have 5,000 square feet. Write an equation that illustrates this situation using the same variables you used in part e.

g) Use substitution to determine the square footage of each house. Demonstrate the process you used to arrive at your answer

Answer by CassieWax(3)

Let the square footage of home #1 be x and home #2 be y.
e) If the square footage of y is 200 more than x, then y=x+200.
f) If the combined square footage of homes x and y is 5,000, then x+y=5,000.
g) From parts (e) and (f), we having the following system of equations:
y=x+200
x+y=5,000
Using substitution to solve, we replace the y in the equation x+y=5,000 with y=x+200
From this, we see that x+(x+200)=5,000.
Solving for x, we see
x+x+200=5,000
2x+200=5,000
2x=5,000-200
2x=4,800
x=2,400
Since x=2,400 and y=x+200, then:
y=2,400+200
y=2,600